OpenStudy (anonymous):

how do you symplify 33/69

OpenStudy (bahrom7893):

divide both top and bottom by 3

OpenStudy (anonymous):

33/69 = (3*11)/(3*23) = 11/23

OpenStudy (anonymous):

Thank you